Then n x n matrix A= [aij] is called a diagonal matrix if aij= 0 when i =/= j. Verify that the product of two n x n diagonal matrices is again a diagonal matrix, and give a simple rule for determining this product.
1
Expert's answer
2020-12-01T05:49:19-0500

let:

"A = \begin{bmatrix}
 a_{11} & 0 & 0 & \ldots & 0 \\
 0 & a_{22} & 0 & \ldots & 0 \\
 \ldots & \ldots & \ldots & \ldots & \ldots \\
 0 & \ldots & 0 & 0 & a_{nn}
\end{bmatrix}"


"B = \begin{bmatrix}
 b_{11} & 0 & 0 & \ldots & 0 \\
 0 & b_{22} & 0 & \ldots & 0 \\
 \ldots & \ldots & \ldots & \ldots & \ldots \\
 0 & \ldots & 0 & 0 & b_{nn}
\end{bmatrix}"


"C = \begin{bmatrix}
 c_{11} & c_{12} & \ldots & c_{1n} \\
 c_{21} & c_{22} & \ldots & c_{2n} \\
 \ldots & \ldots & \ldots & \ldots \\
 c_{n1} & \ldots & c_{n2} & c_{nn}
\end{bmatrix}"


Then:

"c_{11} = a_{11} \cdot b_{11} + 0 \cdot 0 + \ldots + 0 \cdot 0 = a_{11} \cdot b_{11} \\
c_{12} = a_{11} \cdot 0 + 0 \cdot b_{22} + \ldots + 0 \cdot 0 = 0 \\
\ldots \\
c_{1n} = a_{11} \cdot 0 + 0 \cdot 0 + \ldots + 0 \cdot b_{nn}= 0 \\"

"c_{21} = 0 \cdot b_{11} + a_{22} \cdot 0 + \ldots + 0 \cdot 0 = 0 \\
c_{22} = 0 \cdot 0 + a_{22} \cdot b_{22} + 0 \cdot 0 + \ldots + 0 \cdot 0 = a_{22} \cdot b_{22} \\
\ldots \\
c_{2n} = 0 \cdot 0 + a_{22} \cdot 0 + \ldots + 0 \cdot 0 + 0 \cdot b_{nn}= 0 \\"

"\ldots \\
\ldots \\
\ldots \\"

"c_{n1} = 0 \cdot b_{11} + 0 \cdot 0 + \ldots + a_{nn} \cdot 0 = 0 \\
c_{n2} = 0 \cdot 0 + 0 \cdot b_{22} + \ldots + a_{nn} \cdot 0 = 0 \\
\ldots \\
c_{nn} = 0 \cdot 0 + \ldots + 0 \cdot 0 + a_{nn} \cdot b_{nn}= a_{nn} \cdot b_{nn} \\"

It is shown that after multiplication of matrix A and B we receive matrix C which have non-zero diagonal elements, namely: "c_{ij} = 0, i \ne j, c_{ii} \ne 0,i=1,2,\ldots,n" . So C is diagonal matrix.


And formula of multiplication of two diagonal matrices is:

"A \cdot B = C \\
c_{ii} = a_{ii} \cdot b_{ii}, i = \overline{1,n} \\
c_{ij} = 0, i \ne j"